The locus of point of intersection of tangent to parabola y`2=4ax and subtend an angle ¥ ..

Grade:12th

1 Answers

Arun
6 years ago
 
Let m1=m and m2=2m eqn of tangent of parabola=y=mx+a/m from y=4ax^2 a=a y=2m+a/2m eq1 y=mx+a/m eq2nd solving eq we get y=3mx put m=y/3x in eq 2 you get result